OpenStudy (rosajovannaportillo):

simplify (2+i)-(-3+9i)

OpenStudy (zzr0ck3r):

hint: -(3+9i) = -3-9i and i-9i = -8i

OpenStudy (rosajovannaportillo):

i don't know how to solve this. can you show me how?

OpenStudy (sshayer):

\[\left( a+\iota b \right)-\left( c+id \right)=\left( a-c \right)+\iota \left( b-d \right)\]

OpenStudy (rosajovannaportillo):

5-8i?

OpenStudy (rosajovannaportillo):

@zzr0ck3r

OpenStudy (sshayer):

correct.

OpenStudy (rosajovannaportillo):

thank you!!